OpenStudy (anonymous):

Evaluate the expression. -4√-14+50 a.)50 b.)-24 c.)32 d.)-28.3

OpenStudy (anonymous):

I NEED HELP !

myininaya (myininaya):

is that a square root?

OpenStudy (anonymous):

yes . . .

OpenStudy (anonymous):

iqot 32 .

myininaya (myininaya):

so that negative number is inside the square root? none of your answers are imaginary though so I don't think that's what you should had typed here

OpenStudy (anonymous):

yes -14 is inside the square root .

myininaya (myininaya):

then none of the following is a solution

OpenStudy (anonymous):

why whaa did you qet ?

myininaya (myininaya):

if you have a square root of a negative number then it is imaginary and all of the answers you have above are real solutions

OpenStudy (anonymous):

so wouLd it be any answer ?

myininaya (myininaya):

yes but none of the above

myininaya (myininaya):

-4i*sqrt(14)+50

myininaya (myininaya):

are you sure that negative is inside the square root?

OpenStudy (anonymous):

answer is (b) -4√-14+50 = -4 √36 = -4 * 6 = -24